How many ways can five students sit in a row of eight chairs for a photograph?: How many ways can five students sit in a row of eight chairs for a photograph?
Answer by stanbon(18999)

# of ways to pick 3 empty chairs = 8C3 = 8*7*6/1*2*3 = 56
# of ways to arrange the 5 students: 5!
------------------------------------------
Final Answer: 56*5! = 6720 ways
----------------------------
Cheers,
Stan H.